I'm supposed to find the specific term of each arithmetic sequence. The problem is : t1=5, t3=20; and I need to find t subscript 12. I can use the equation tn=t1+(n-1)d if that helps you any? Thanks for the help.
Answer by htmentor(1343)

The n-th term of the arithmetic sequence is written t_n = t1 + (n-1)d where
d = the common difference and t1 is the 1st term
So we need to find d
We can form two equations for the common difference
d = t2 - t1 = t2 - 5
d = t3 - t2 = 20 - t2
Add the two equations together:
2d = 15 -> d = 7.5
Therefore, t12 = 5 + 7.5(12-1)
t12 = 87.5